Solve for x and use radicals as needed:
x^2 + 18x + 4 = 0
Thank you!!

Answer by stanbon(75887)

Solve for x and use radicals as needed:
x^2 + 18x + 4 = 0
Solve for x and use radicals as needed:
x^2 + 18x + 4 = 0
---
x = [-18 +- sqrt(18^2 - 4*1*4)]/2
============
x = [-18 +- sqrt(48)]/2
---
x = [-18 +- 4sqrt(3)]/2
---
x = [-9 +- 2sqrt(3)]
---
x = -9+2sqrt(3) or x = -9-2sqrt(3)
===================================
Cheers,
Stan H.
==========